OpenStudy (anonymous):

1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? (2 points)

terenzreignz (terenzreignz):

Okay, Let x = Tony's age. then Belinda's age is: 2x + 8 ----> since Belinda's age is 8 more than twice Tony's now, their combined age is is 56, so: x (Tony's age) + 2x + 8 (Belinda's age) =56 x + 2x + 8 = 56 3x = 56 - 8 3x = 48 x = 16 So Tony's age is 16 Belinda's age is 40 (twice 16 plus 8)